### v1

# The Higgs Mass and a Perspective on Broken Autonomy of Scales

Abstract.
The hierarchy problem of the Higgs mass, the violation of the autonomy of far-distant energy scales, is identified as a pseudo-problem. The pseudo-problem is based on the dogmatic adherence to the methodology of effective theories, for which there is no justification when dealing with presumably fundamental phenomena such as the Higgs mechanism. In view of further breaches of autonomy of scales in fundamental principles outside particle physics, the hierarchy problem is instead reinterpreted as an indication of the fundamental ontological status of the Higgs boson. Selective Realism justifies this attribution within the effective theories of the standard model of particle physics. Moreover, a model of changes in aesthetic canons provides an explanation for the existing dogmatism regarding effective theories.
Key words and phrases:
Physics Beyond the Standard Model, Naturalness, Autonomy of Scales, Hierarchy Problem, Effective Theories, Reductionism
1. Introduction
This article deals with the naturalness problem of the Higgs mass, also known as the hierarchy problem. Despite numerous attempts to characterize the problem, to question its status as a problem, or to address it with new physics: One is far from a consensus, especially since, in some interpretations, it questions the theoretical framework of particle physics and has profound philosophical implications. In particular, the past decade has seen a fruitful exchange between physicists and philosophers of science, which this work seeks to join.
The naturalness principle, independent of its precise definition, regards the physicist’s expectation of numerical values for parameters. By a majority that still shapes the research program, theories in particle physics are still considered unnatural and in need of explanations, if such numerical values turn out to be particularly small or large in experiments. Naturalness is percepted as the central concept "if one has to summarize in one word what drove the efforts in physics beyond the standard model of the last several decades" (Giudice 2017, p. 3). Naturalness has become a core criterion in theory assessment in the standard model of particle physics as well as physics beyond the standard model (BSM physics for short), and its violation has been and is still considered as something in need for an explanation. It may seem surprising that the efforts of BSM physics have been significantly driven for half a century by such an inconspicuous criterion as the order of magnitude of certain constants of nature. Hence, given that significant progress in the history of physics has been made due to unexplainable empirical observations or theoretical inconsistencies (which both do not occur in case of unnatural parameters), naturalness requires a critical revision of its validity and relevance as a criterion for theory evaluation. The particular urgency of a revision arises from the previous empirical failures of all attempts to explain unnatural parameters using new physics.
An upfront concession: In the virtual absence of unexplainable empirical results in the BSM physics research program (except of neutrino oscillations and possibly dark matter) with, simultaneously, some plausibility for BSM physics at the yet unknown energy scales up to the Planck energy, resorting to guiding principles like naturalness seems initially legitimate. In a highly developed research program where fundamental limits of the technically possible are imposed in the search for empirical evidence, astonishment about certain numerical values should also be paid attention to. This is especially the case when unnatural values impact two cornerstones of theory-building in this research program: symmetries and effective field theory, namely in the definitions of ’t Hooft naturalness (1979) and technical naturalness (Susskind 1979). These concepts attribute unnaturalness to the $\bar{\theta}$ parameter of QCD (strong CP problem) and the Higgs mass in the standard model. Both might go beyond mere aesthetic discomfort. However, numerous attempts to solve both problems with new physics never led to empirical progress. Additionally, the debate on technical naturalness becomes even more significant with its interdisciplinary emergence, as it also arises in the context of the cosmological constant [^1] See (Craig 2022) for an overview of the strong CP problem, the hierarchy problem and the cosmological constant..
This article claims that the restoration of technical naturalness of the Higgs mass with new physics represents an engagement with a pseudo-problem, thus hindering the progress of the BSM physics program. We will present arguments that there is no need for a physical explanation of a violation of technical naturalness in this particular case. This far-reaching critique of the presumably most crucial non-empirical guiding principle of BSM physics, which significantly contributed to the multibillion-dollar construction of the LHC in Geneva [^2] Of course, confirming the hypothetical Higgs boson itself was perhaps an even stronger reason to build the LHC. However, this discovery was not its only purpose., is not isolated. In recent times, several calls for a re-evaluation of naturalness have been made. Prominent critics include the CERN theorist Giudice (2017) and the author, physicist, and philosopher Hossenfelder (among others: 2018, 2019). While Giudice, formerly a staunch defender of naturalness, offers mild criticism and proclaims a post-naturalness era but wishes to retain naturalness as a still significant criterion, Hossenfelder is uncompromising. She calls for a complete departure from the principle of naturalness, which she believes has misled particle physics for too long [^3] Other contributions to the naturalness debate include (Giudice 2008, Wells 2013 Dine 2015, Hossenfelder 2017, Williams 2015, 2019, Wallace 2019, Bain 2019, Franklin 2020, Fischer 2020, Koren 2020, Athanasiou 2022). Our apologies go to all authors whose contributions were not mentioned..
In this article, at least some significance should be attributed to the value of the Higgs mass, not creating a need for explanation, but suggesting that the Higgs boson is ontologically fundamental and placing it outside the provisional, effective concepts of the standard model physics. The breakdown of autonomy of scales marks the end of an instrumentalist, pragmatic approach in particle physics that cannot meet the claim of fundamental physics to make definitive statements about the world - ultimately, the desire to relieve philosophy of the task of ontology. In short, instead of being horrified by the failure of effective theories, a sensitivity to large energy scales should provide ample reason for optimism that the overarching goal of fundamental physics will succeed.
The argumentation will proceed as follows: In Chapter 2, we first recapitulate some important definitions of naturalness and particularly describe the lack of technical naturalness of the renormalized Higgs mass. Confronted with this problem, there are three options: firstly, attempting to address it with new physics but repeatedly facing disappointment due to the experimental failures of all previous theories. Secondly, attempting to completely deny the problem: the Higgs mass is a constant of nature, and discussing its value becomes unnecessary. We opt for the third option and question the adherence to effective field theories, but derive a positive reaction to the violated autonomy of scales. To do so, we argue in Chapter 3 that quantum corrections to the Higgs mass can indeed be ascribed a physical reality, and the dependence of the Higgs mass on high-energy domains has significance for theory evaluation: the breakdown of autonomy of scales. However, we then narrow down the scope of autonomy of scales by explaining several phenomena in which its violation occurs. For each of these phenomena, a fundamental ontological status can be attributed, and the violation of autonomy of scales does not pose a problem. In Chapter 4, we argue against the dogmatic adherence to this principle and thus to effective theories as the only way to do physics, as the philosophical consequences of this adherence are incoherent with the actual modus operandi of physics and the worldview revealed therein. Finally, with Selective Realism, we attempt to reconcile the fundamental nature of the Higgs mass with the effective theories of the standard model. Chapter 5 aims to provide the philosophical framework for (irrational) dogmatic adherence to autonomy of scales through McAllister’s theory of scientific change by the change of aesthetic ideals. In the conclusion of Chapter 6, the significant role of effective theories for a thriving future particle physics is emphasized. Nonetheless, occasional detachment from their principles is necessary to avoid pseudo-problems like that of the unnatural Higgs mass.
2. Technical Naturalness and the Hierarchy Problem
2.1. Definitions of Naturalness
In physics in general, the concept of naturalness arises in the context of Dirac’s Large Number Hypothesis: Given an operator $\xi$ of the form $\mathcal{L}\propto c_{0}\xi$, where $\mathcal{L}$ represents the Lagrangian of the physical theory, the dimensionless parameter $c_{0}$ in natural units is expected to be of the order $\mathcal{O}(1)$ (Dirac 1974). Throughout the last century, Dirac repeatedly appears as an advocate of aesthetic criteria as a decisive basis for evaluation in physics. Similarly, this notion does not have underlying physical principles justifying a dogmatic adherence to absolute naturalness (cf. Wells 2013). It appears to be solely based on a desire for numerical aesthetics. Although the principle of absolute naturalness often proves correct, it fails for many parameters of the standard model and must, therefore, be considered too broad and unsuitable for further discussion.
In the context of the standard model, to the best of my knowledge, Weinberg (1972) first mentions naturalness in an attempt to find a "natural explanation of the approximate symmetries in nature". Symmetry, whose profound significance for particle physics is undoubted, can restrict the demand for absolute naturalness. This ’t Hooft naturalness (1979) allows $c_{0}$ to take a much smaller value than $\mathcal{O}(1)$ as long as the limit $c_{0}\to 0$ leads to a symmetry. Chiral symmetries or gauge symmetries, for example, protect the fermion masses or gauge boson masses of the standard model from (unnatural) large correction terms in interaction with virtual particles, justifying the constraint imposed by ’t Hooft. Mathematically, this cannot happen for the Higgs boson since this particle, being the only one, has no spin - its mass is ’t Hooft-unnatural. The definition of naturalness that is more relevant for this chapter is reproduced from the ’t Hooftian perspective but further narrows down the set of unnatural phenomena. In this work, this is defined as technical naturalness [^4] In the literature, this term is sometimes used synonymously with ’t Hooft naturalness. Instead, in the present context, the term AoS naturalness (autonomy of scales) is used. However, the unfortunate term technical naturalness, sounding somewhat contradictory, cannot be avoided even in this convention.. It describes the prohibition of mutual influence from energy scales separated by many orders of magnitude (autonomy of scales) and first appears in Susskind (1979). We will sharpen that definition after having provided somewhat more background.
Symmetries, in the spirit of ’t Hooft naturalness, can preserve technical naturalness, as seen previously with the masses of the SM. However, the small $\bar{\theta}$ parameter of QCD does not violate autonomy of scales but is ’t-Hooft-unnatural. For this reason, the violation of technical naturalness must already be considered a clearer indication of a need for explanation, as it implies ’t-Hooft unnaturalness (the absence of a protective symmetry against a correspondence between widely distant scales) - at least if naturalness is considered a valid guide in general. Furthermore, the demand for autonomy of scales is closely related to the construction principle of an effective field theory (EFT), which, in the standard model, uses the cutoff parameter in the renormalization of particle masses. In contrast, the desire for approximate symmetries when finding small parameters according to ’t Hooft appears more as a straightening of the aesthetics of the theory: If one must accept unsightly parameters, they should at least come with a symmetry. Criticism of the present modus operandi of BSM physics, accessible to a broader readership in content, occasionally suggests that physicists are more guided by aesthetic requirements for parameters than the genuinely physically, substantial definition of technical naturalness reveals. This is not to deny the actual importance of aesthetics in the search for BSM physics but to clarify the theory-internal significance of technical naturalness and separate it from the aesthetics of science. It sometimes conveys a misleading impression in the scientific literature when both the strong CP problem and the hierarchy problem of the Higgs mass are mentioned in one sentence under the umbrella term of naturalness (see, for example, Hossenfelder 2018, Giudice 2017, Craig 2022). These are two very different concepts.
2.2. Renormalization of the Higgs Mass
Determining the absolute strength of fundamental interactions in the Universe, one is confronted with a very clear hierarchy (giving the naturalness problem its name, as explained now). The extremely weak effect of gravity compared to all other forces, with a factor of $10^{-41}$ compared to the strength of electromagnetic interaction on the quark length scale, is initially fortunate: Gravity only acts as a dominant interaction in the case of large mass accumulations (such as planets and stars), whereas in the subatomic realm its effect can be neglected and no quantum gravity is needed yet.
Unfortunately, gravity returns through the back door into quantum field theory as we venture closer to the Big Bang, involving higher temperatures of the universe, and thus higher particle energies. On this journey, one traverses "the desert", the unknown physics that can extend over 15 orders of magnitude up to the Planck scale. Due to the lack of a theory of quantum gravity, precise statements about processes at the maximum of this scale cannot be made. Therefore, in quantum field theory, each integral over particle energies is cut off where new physics is expected, and quantum field theory is superseded - with the cutoff $\Lambda\propto M_{\mathrm{Planck}}$ (or already at $M_{GUT}$). The standard model becomes an EFT.
The procedure of renormalization removes the cutoff-dependence: In the initial formulation of renormalization during the development of QED, the limit $\Lambda\to\infty$ was considered. This formalistic idea of renormalization, which caused discomfort to physicists of this generation precisely because of its infinities (cf. Kragh 1990, p. 1984 on Dirac), fails due to the fundamental limits of the applicability of QFT (quantum gravity, Landau poles). Wilson’s (1975) significant contribution, a representative of the next generation of quantum field theorists, was ultimately based on the reinterpretation of formalistic renormalization as realistic renormalization with a finite cutoff. He conceives QFT as a limited, effective field theory and exemplifies this idea with the requirement $\Lambda\propto M_{\mathrm{Planck}}$. This contemporary understanding of renormalization reveals that QFT cannot be the ultimate answer to questions in fundamental physics. However, the finite cutoff of the renormalizable theory is finally the reason "that we can do sensible macroscopic physics even without having detailed knowledge of the underlying microscopic theories" (Adlam 2019, p. 10).
With these preparatory remarks, we can turn to the Higgs mass: After renormalization, the parameter of the bare Higgs mass $m_{H,b}$ is replaced by the renormalized Higgs mass $m_{H,r}$, into which the perturbative corrections (self-energy) $\Delta m_{H}$ went. More precisely: $m_{H,r}=m_{H,b}+\Delta m_{H}$. The calculation of fermionic loops in first order yields the following correction:
$\Delta m_{H}^{2}=\frac{\Lambda^{2}}{16\pi^{2}}\biggl{(}-6y_{t}^{2}+\frac{9}{4}g^{2}+\frac{3}{4}g^{\prime 2}+6\lambda\biggl{)}$
The exact parameters (Yukawa coupling to the top quark $y_{t}$ as well as parameters from the Higgs potential and electroweak symmetry breaking) are less important than the huge prefactor $\Lambda^{2}$, which represents the upper limit of the theory’s validity - worst case $10^{18}$ GeV (so-called UV range of the theory). The experimentally extracted renormalized mass $m_{H,r}$ of about $125$ GeV (so-called IR range of the theory) requires absorbing up to 16 orders of magnitude in the bare mass. The relatively tiny remaining value is perceived as unnaturally small and thus in need of explanation, as the widely separated energy scales must be artificially separated through the choice of the bare mass (not measurable). This does not happen for all other elementary particles because their loop corrections only contain terms of the form $\log(\Lambda)$ and thus limit their order of magnitude to $\mathcal{O}(10^{1})$.
Moreover the following must be noted: UV sensitivity of the Higgs mass is only present in the sense that large terms from the UV sector appear in the corrections. This fact must be given attention as it represents at least a flaw for the concept of effective field theories (EFT). However, it does not mean that the Higgs mass could be sensitive to small changes in UV parameters (as would be feasible, for example, in changing initial conditions in dynamic systems) - these values are immutable.
Nevertheless, the problem of technical naturalness is not sharply separated from problems of accessing the world through effective theories; there is mutual influence. In recent literature, there has been a debate about the precise definition of technical naturalness and its influence on theory-building using EFTs. Giudice (2017, p. 3) initially stated: "Naturalness is an inescapable consequence of the ingredients generally used to construct effective field theories." Hossenfelder (2017) then replied: "Of course it is not. If it was, why make it an additional requirement?" In a strict interpretation of Giudice’s claim, an error can indeed be identified. Requiring technical naturalness is not equivalent to the requirement of decoupling energy scales, indispensable for the successful application of an EFT. Nevertheless, its violation raises suspicions about the universal applicability of EFTs. Williams (2019, p. 17) attempts to reconcile by admitting: "(Technically) unnatural theories seem to violate the spirit, though not the letter, of the relationship between scales in effective field theory." Technical naturalness is thus a stricter principle. An EFT does not initially provide information about how large the quantum corrections $\Delta m_{H}$ is allowed to be. Values from the UV range do not yet contradict the use of EFTs. He explains: "Naturalness in this sense is motivated insofar as structural features of effective field theory lead us to expect a general insensitivity of low-energy observables to the structure of the theory at much higher energies, but it goes beyond what is strictly licensed by those structural features" (Williams 2019, ibid.).
A clear distinction decoupling and technical naturalness has already been attempted. Franklin (2020) differentiates between autonomy of microstates (IR dynamics is autonomous despite changes in the UV range) and autonomy of microlaws (IR dynamics is autonomous despite changes in UV parameters) [^5] In a very similar way, Bain (2019) calls this difference heuristic and precise decoupling.. However, for the lessons learned from the subsequent argumentation, it will be more helpful to differentiate between weak and strong autonomy of scales: Weak autonomy guarantees that an effective theory $T^{eff}_{i}$, defined on an interval $[L_{i},\Lambda_{i}]$ (IR cutoff $L_{i}$ and UV cutoff $\Lambda_{i}$) empirically adequately describes its intended phenomena, i.e., on intervals $[0,L_{i}]$ and especially $[\Lambda_{i},\infty)$, there are no crucial additional effects that could seriously influence the behavior of the phenomena. In other words, it concerns the real correspondence to distant scales. This weak autonomy legitimizes the application of EFT. In the case of strong autonomy, contributions to phenomena in $T^{eff}_{i}$, expressed in units of $\Lambda_{i}$, are roughly of the order of $\mathcal{O}(10^{1})$. However, their violation does not prevent the application of EFT due to the option of renormalization. The quantum corrections are only a crucial additional effect in the non-renormalized equations, which is then mathematically manageable. In other words, it concerns the virtual correspondence to distant scales.
An analogy may help: A researcher is confronted with a comprehensive textbook on the knowledge of an entire discipline. They want to learn everything important about a specific topic in this discipline and read the book effectively: In the table of contents they find that the topic is covered in Chapter 3, so they focus on this section, and ignore the rest of the book. Weak autonomy is violated if crucial information on the subject can be found in Chapter 28, which cannot be neglected to understand it. The researcher would now have to make significant efforts and read the entire book. Strong autonomy is already violated if, for example, a footnote in Chapter 3 refers to additional content in Chapter 28. Although the content is artificially brought into the limited section of Chapter 3 via the footnote, and effective reading is not hindered, suspicion arises: The topic could have greater importance than initially assumed, could possibly occur universally and fundamentally, if there are correspondences to completely different sections of the book.
With this illustration, it shall be clarified that even strong autonomy of scales is always closely intertwined with the application of effective theories. Both are expressions of the same expectation to the world: that there should be no correspondence between UV and IR ranges. The violation of strong autonomy already questions whether effective theories provide the appropriate framework to adequately represent all phenomena occurring in them. In the case of the technically unnatural Higgs mass, an EFT remains usable; however, it takes this parameter as an artifact that would presumably need to be assigned to a different type of theories (fundamental). For this reason, aspects related to both strong and weak scale autonomy will be discussed in the following, and the critical examination of adherence to effective theories is inseparably linked to the revision of the hierarchy problem, too.
2.3. How to Not Solve the Problem
Several arguments cannot resolve this naturalness problem, as has already been highlighted by other authors:
  • The calculation of corrections in dimensional regularization makes the cutoff disappear. Yes, but the corrections will also take a huge value, as now the heaviest fermion mass $M^{2}$ of the effective theory takes its role (Williams 2015).
  • Quantum gravity might not impose a limit on the Higgs boson, but be valid in isolation. The limit $\Lambda\to\infty$ would be justified, and the Higgs mass would be natural. Yes, but the gauge coupling to hypercharge reaches its Landau pole around $10^{41}$ GeV. The theory remains effective (Craig 2022), unless progress in rigorous QFT solves the general problem of appearing Landau poles and provides a changed perspective on the hierarchy problem.
  • Supersymmetry provides scalar loop corrections that could restore naturalness. Yes, but the crucial regions of the parameter space are already excluded. Unexcluded values for loop corrections may lower the value of the corrections, but the theory remains unnatural to some degree (Giudice 2017).
Any argument based on the decoupling theorem (Appelquist, Carazzone 1975) must also be excluded. As expressed in several works (cf. Hartmann 2001), the proof relies on an outdated renormalization scheme and the assumption of the general renormalizability of theories, which has already been ruled out for quantum gravity.
Finally, let’s consider the discussion around probability distributions of parameters diagnosed with unnaturalness. The absence of naturalness is often perceived as a problem because small quantities that do not allow variation seem improbable. As emphasized regularly in the literature (Hossenfelder 2019, Williams 2019), however, such a statement requires a probability distribution, which typically implies a multiverse, where, according to that distribution, all possible parameter values are realized in different universes. Otherwise, the attribution of "improbable" would be an unphysical statement, nothing more than a disturbance of aesthetic perception without consequences for the validity of the theory. Due to the ignorance about the shape of such a distribution, if it exists, the demand for naturalness should be abandoned.
Including a probability distribution and the multiverse in the discussion about naturalness seems, however, unhelpful. Firstly, from a pragmatic viewpoint, the naturalness debate does not gain higher physical content with the addition of a multiverse. The concept of a multiverse in the sense of string theory [^6] The multiverse needed for the argumentation here is clearly distinguished from the multiverse of Everett’s many-worlds interpretation of quantum mechanics (Everett 1957). In Tegmark’s (2008) definition, we are dealing with a Level II multiverse in this case. remains a purely mathematical concept that evades any empirical verification and presumably always will. But let’s assume that one could derive a probability distribution of the relevant parameter values and empirically test it. Shifting the discussion about specific values in nature into the multiverse would then represent a complicating relocation of the problem without obtaining insight, as we shall argue in the following. No matter how probable the discovery of values in our universe is, the likelihood is not zero by construction of the distribution. After all, the values are realized in at least one universe, being ours. Turning the naturalness debate into a discussion of a probability distribution of parameters in the multiverse is intended to answer the question of whether we live in a probable realization of SM parameters - hence a natural world without the need for explanation. However, due to the concrete realization of any parameter value according to a given distribution in some universe, the need for explanation is equally absent in every conceivable set of parameters. Noticing that our universe represents an improbable parameter constellation would not be a problem but an unassessable fact. Knowledge of a distribution is therefore irrelevant. With or without the multiverse, the tautological statement remains: The parameters are what they are. It is inconsequential whether they are unchangeable (universe) or variable and possibly improbable (multiverse) - in both options, they cannot be called unnatural.
3. Autonomy of Scales in Effective and Fundamental Theories
Confronted with the experimental value of the Higgs mass, one initially has several options to respond to the dependence of perturbative corrections on physics at extremely high energies: The BSM enthusiast suspects new physics and is disappointed by the experiments of the past decades (Giudice 2017). I’m not saying these experiments were in vain. If a single SM particle evades the otherwise reliable decoupling of low-energy scales from UV physics, this may have an explanation within particle physics. Temporarily ruling this out can and should be considered asa non-negotiable part of scientific practice. The disappointed one can now simplify things and claim: Only the measured physical value of the Higgs mass has significance for the theory, a constant of nature: It is what it is, and any discussion about its numerical value is unnecessary (like Hossenfelder 2018). This would be a prime example of a pseudo-problem. However, as we will see below, this other extreme cannot be an appropriate answer. So, the third option remains to break out of internal physical reasoning and question the methodology of theoretical particle physics (the exclusive use of effective theories) itself. This has already happened in the recent past, albeit with sometimes drastic consequences for physics. Wallace (2019), for instance, warns that the rejection of the principle of naturalness implies a rejection of reductionism as a overarching principle of foundational research. In the following, however, a positive consequence of rejecting the principle of naturalness will be argued for.
3.1. Bare Masses and Loop Corrections
Firstly, it is decisive to clarify the physical reality of the various elements of the renormalization process in order to argue against the second option. In characterizing the hierarchy problem, Wetterich (1984, p. 217) writes in one of the earliest analyses: "In any case, fine-tuning of bare parameters is not really the relevant problem (…) The relevant parameters are the physical parameters, since any predictions of a model must finally be expressed in terms of these". It is, therefore, unproblematic to compensate for the enormous correction terms $\propto\Lambda^{2}$ by an appropriate bare mass; in the end, the renormalized mass remains as the only isolated measurable quantity. Whether the correction terms have a dependence $\propto\log(\Lambda)$ or $\propto\Lambda^{2}$, there is no problematic fine-tuning, and the hierarchy problem loses significance. This part of Wetterich’s analysis of the hierarchy problem is undoubtedly correct.
Indeed, the bare mass should not be regarded as an intrinsic property, as we will argue now. From a philosophical viewpoint, one could argue against considering objects with bare mass as independent entities as well: Speculating about the significance of its numerical value would be akin to trying to quantify the charge of a lone electron in an otherwise empty hypothetical universe. Without a role in a relational framework (interaction), this mere information is meaningless. Ontologically spoken, this argument can be described most accurately with the non-eliminative version of ontic structural realism (OSR) à la Chakravartty (2012). This variant of the original OSR, which itself goes back to Ladyman (1998) in its more drastic eliminative formulation, limits the ontological status of irreducible objects to existing as individual relata of a relation, so that the relata possess only relational and no intrinsic properties [^7] In contrast, eliminative OSR denies the existence of objects carrying intrinsic properties altogether; only physical structures should exist.. Applying this to renormalization: Physical particle masses are inevitably dependent on interactions with their environment, specifically on the fluctuations of the vacuum in the form of virtual particles. The vacuum, with its ability to influence the bare mass (and likewise the bare charge), is part of physical reality and not an independent stage on which the course of the world takes place. Embedded in the structure of the vacuum (the relation that triggers the interaction of the bare particle with virtual particles), the elementary particle (relatum) consequently acquires its renormalized mass (and possibly charge) as relational properties. The existence of bare mass as an intrinsic property of the elementary particle is thus denied, and its numerical value is irrelevant, with its fine-tuning being outside the debate.
However, Wetterich’s analysis overlooks the higlighted physical reality of quantum field theoretical corrections in $\Delta m_{H}$ the relation that actually exists in the OSR interpretation [^8] They are as real as the interaction of a macroscopic object with one or another medium: The renormalized (i.e., effective) mass of a ping pong ball crucially depends on whether it is in a water bath, in the air, or in a vacuum. As a composite object, of course, the ping pong ball additionally has an intrinsic bare mass.. For this reason, prematurely dismissing the hierarchy problem as a misguided debate about fixed constants of nature as a mere measurement result is too simple a way out. The unique position of the Higgs boson as the only spin-0 particle in the described relation, which grants it a glimpse into UV physics, has a physical significance: The relation between spin-0 of the real particle and the virtual particles of the vacuum gives rise to a physical phenomenon that occurs in several cases of fundamental physics - the violation of autonomy of scales. Initially, it does not matter at which energy scale new physics emerges beyond the effective theory. In the case of scalar elementary particles, their limit represented by the cutoff $\Lambda$ always leads to a dialogue between the IR and UV regions of the theory, somewhat serving as a link to more fundamental energy scales.
3.2. The Applicability of Technical Naturalness
With these foundations, we can turn to a critical examination of the scope of autonomy of scales in fundamental physics. Consider the impact of the UV-IR dialogue occurring in the context of the Higgs mass on the widely adopted methodology of EFTs. In the tower of quasi-autonomous domains in the sense of Cao and Schweber (1993), each described by an EFT, the existence of a scalar elementary particle breaks this autonomy and establishes a dialogue between significantly different energy scales. Is it a coincidence that this phenomenon appears in several mechanisms in the universe, which can be considered fundamental for good reasons?
In three mechanisms, such a delicate dependence of large length scales or small energy scales on the other end of the spectrum occurs. Firstly, the specific properties of the quantum vacuum at the smallest length scales influence the expansion rate of the entire universe (vacuum energy), which undoubtedly plays out on the largest available length scales when calculating the cosmological constant (see also Craig 2022 for an overview). Unlike the balancing act performed between the electroweak scale ($10^{-18}$ m) and the Planck scale ($10^{-35}$ m) in the hierarchy problem of the Higgs mass, quantum effects in cosmology are responsible for considerably larger length scales. On the other hand, effective theories usually separate cosmology and QFT999 This passage only addressed the principal influence of the quantum vacuum on the largest possible length scales. Moreover, the violated strong autonomy of scales exists in a way mathematically analogous to the Higgs mass: Summing the bare cosmological constant from general relativity and the value of the vacuum energy, $\tilde{\lambda}+\lambda=\Lambda$, results in handling a mismatch of 120 orders of magnitude between quantum effects and the observable value $\Lambda$ (to be distinguished from the cutoff)..
A second UV-IR mixture arises under the assumption made for a theory of quantum gravity that space at the Planck length must have a non-commutative nature. This supplements the theory with the commutator relations between different operators, such as position and momentum, already necessary in quantum mechanics. To prevent gravitational collapse when localizing events with extreme precision ($l_{\mathrm{Planck}}$), one demands uncertainty conditions $\Delta x^{\mu}$ in the form $\Delta x^{0}(\Delta x^{1}+\Delta x^{2}+\Delta x^{3})\geq l^{2}_{\mathrm{Planck}}$ and $\Delta x^{1}\Delta x^{2}+\Delta x^{2}\Delta x^{3}+\Delta x^{3}\Delta x^{1}\geq l^{2}_{\mathrm{Planck}}$ in four dimensions. These uncertainties are induced by the commutator $[x^{\mu},x^{\nu}]_{-}=i\theta^{\mu\nu}$ (Doplicher, Fredenhagen, Roberts 1995). In prose: If one wants to determine a coordinate with maximum precision (UV), a maximum uncertainty (IR) remains for the value of the other coordinate. This mechanism will probably represent a fundamental property of spacetime when quantum effects of gravity need to be considered.
UV-IR mixtures thus appear as a generic feature of non-commutative geometries, as pointed out by Minwalla, van Raamsdonk, and Seiberg (2000). Effective theories of gravity, as well as QFT, use the commutative limit $\theta\to 0$, in which this mixture does not occur. Koren (2020, p. 265) illustrates as follows: "An effective field theorist living in a noncommutative space would have no way to understand the appearance of this infrared scale; its existence is intrinsically linked to the geometry of spacetime and to the far UV of the theory". Similarly, an EFT euphoric deals with the Higgs mass, but attempts to conceal a probably inherent fact with new physics. Surprisingly, UV-IR mixing is readily accepted when constructing a non-commutative spacetime.
A third sensitivity arises already in the mesocosm itself, without the need to know anything about quantum effects: in the theory of complex systems. This represents a significant phenomenological extension of linear Newtonian physics without leaving classical physics. In linear mechanics, the dynamics of a system is unaffected by infinitesimal disturbances. The dynamics of nonlinear systems, on the other hand, show a sensitive dependence on the initial conditions with the slightest influence - small changes have a significant effect (UV-IR mixing). This is particularly evident in the butterfly effect, provocatively expressed in the question of whether the flapping of a butterfly’s wings in Brazil can trigger a tornado in Texas (Lorenz 1972). More precisely, the presence of turbulence in a multitude of complex systems, extremely dependent on tiny perturbations of initial conditions, hints that a manifestation of broken autonomy of scales occurs as a fundamental property in the mesoscopic world as well, while the linear, i.e. effective theory describes only a small fraction of deterministic phenomena that show no dependencies on distant energy scales.
Contrasted with violations of autonomy of scales, I refer to three successes of technical naturalness in quantum field theory: First, the divergent electron self-energy in classical electrodynamic, cured by the postulate of positrons by Dirac. Second, the pion mass difference, which can only be called technically natural, if the cutoff $\Lambda$ of its effective theory appears around 800 MeV, cured by the later discovered $\rho$ meson. Third, the mass difference between the long- and short-lived state of the $K$ meson (kaon). Empirical reasons demand $\Lambda<3$ GeV, which led to the prediction of the mass of the charm-quark [^10] Instead of citing all single publications, we refer to (Williams 2015), where all three phenomena are explained in greater detail.. All three successes of technical naturalness share the commonality that the phenomena to which the principle was applied must not be considered fundamental, in stark contrast to the three cases previously discussed. Instead, they each occur in effective field theories (self-explanatory for classical electrodynamics) and involve composites (mesons) of elementary particles described in a narrow interval of the energy spectrum. It is helpful to remark, as Williams (2015) has already outlined, that only the last phenomenon was explained by explicitly applying the demand for naturalness. The former ones were recognized retrospectively as confirmations of the naturalness principle in QFT.
We do not know why the decoupling of distant energy scales is a widespread phenomenon that simplifies life for physicists. In addition to the mentioned cases from QFT, it holds equally true in classical physics. It is the reason why physics could gradually approach more challenging theories, and the description of an apple falling from a tree did not require quantum gravity. It is the conditio sine qua non of an effective theory. However, there is no reason to assume that (both weak and strong) autonomy of scales must be preserved at the end of the chain of effective theories. Quite the opposite: An effective theory aims to provisionally and simplifiedly represent a part of reality (a section of the entire energy scale) - a fundamental theory, on the other hand, captures elements of reality itself (not limited to specific energy ranges). A violation of autonomy of scales, in both definitions, seems almost a natural consequence [^11] It should be noted that this viewpoint is far from consensus: Crowther (2019) provides a list of criteria a fundamental theory shall fulfil and explicitly mentions naturalness as one criterion..
4. Beyond Effective Field Theories
The absence of technical naturalness becomes a cause for concern only under the assumption being discussed: that the seemingly reductionist theory sequence is endless and the series of effective theories does not terminate. Lacking autonomy of scales would indeed signal the failure of reductionism. In this case, against whose plausibility we will now present arguments, the general modus operandi of foundational physics would be challenged to such an extent that it must be elevated to the status of a naturalness problem, rather than characterizing it as an almost desirable trait. We will see below how influential publications have cemented the adherence to EFTs, thereby hindering a comprehensive consideration of the hierarchy problem.
4.1. The World as a Collection of Quasi-Autonomous Domains?
In the following discussion, we revisit the ontological consequences of considering autonomy of scales. To strengthen the argument for the reevaluation of the absence of technical naturalness, we turn to a widely discussed work by Cao and Schweber (1993), which describes physics as the aforementioned infinite tower of quasi-autonomous domains, each ontologically equivalent and effectively described. Thus, the authors not only adopt an anti-reductionist and anti-fundamentalist stance (infinite tower), but they also postulate an ontological pluralism, thereby inadequately representing the methods and understanding of physics in the last century, as we will see below. While in the first part of the work, they meticulously trace the history of renormalization, they derive significant consequences from Wilson’s realistic reinterpretation of formal renormalization into effective field theory (1975) in the second part. These consequences have been criticized extensively (e.g., Hartmann 2001), partly due to the use of the decoupling theorem (Appelquist, Carazzone 1975), which was excluded from the argument list for good reasons beforehand. Nevertheless, I will once again challenge the worldview outlined there, as the scope of the work has influenced the assessment of absent autonomy of scales.
It is helpful to recap the mode of operation of a foundational physicist: She deems a theory effective because she understands it as a provisional approximation to a cosmos governed by a fundamental theory. At the time of developing the EFT, this fundamental theory is evidently not within intellectual and perhaps empirical reach yet, so one settles for an approximation reliable in certain domains. During the convergence (in its empirical adequacy) of the sequence of theories reliable in certain domains against a fundamental theory reliable in all domains, it represents a precursor of merely epistemic value. Instead of making ontologically fundamental statements already, it simply propels the physicist a significant step forward on this path. This represents a central assumption of foundational physics, abandoning it would instead deny the epistemic value of such interim successes. This pragmatic approach to the world of the smallest, whose exploration is solely driven by the pure desire for knowledge and cannot be motivated by practical significance, is legitimized only by the attainability of a foundation. EFTs thus serve as instruments to approach fundamental theories. This approximation is missing in the infinitely layered tower of EFTs. The potential senselessness in the quest for fundamental physics in this picture is not an argument in itself that can be used against Cao and Schweber. It is rather the fact that ever higher precision and ever smaller distances to the extreme scales of physics (Planck scale) do not give any reason to believe that we are just as far from the foundation of the tower as in Newtonian mechanics, as suggested by the image of the infinite tower. In short, Cao and Schweber paint a false picture of the way of working of a foundational physicist in exploring the world, as well as of the world itself.
Critique can also be directed at other theses: The coexistence of different EFTs by no means allows the conclusion that our world is ontologically plural. An effective theory, by definition, cannot enable ontological determinations. Moreover, no actual autonomy of these respective domains can be established. There are continuous transitions in, for example, the theory of strong nuclear force, where the approximate validity of point-like nucleons is steadily reduced, and at no discrete point does the absolute necessity of quarks become apparent. Such smooth transitions are already present in the construction of the theory to be effective. The transition to another domain is always at the discretion of the physicist, who decides what deviations are tolerable for the given context.
With these arguments, the long-term significance of EFTs and thus the adherence to technical naturalness is once again questioned. Williams (2015) initially concludes that its violation in the case of the Higgs mass is a counterargument to the worldview previously presented by Cao and Schweber (1993). However, even its general consideration (without knowing anything about autonomy of scales) reveals, as shown, sufficient incoherence with the discipline of physics to reject it.
4.2. Selective Realism in the Realm of EFTs
Athanasiou (2022, p. 115) categorizes the work of Cao and Schweber (1993) among that of EFT enthusiasts (among optimists, cautious and sceptics). In contrast, when addressing EFT skeptics, he states, "Clearly, the skeptical road is the path of the patient: in the absence of a completely satisfactory theory, thou shalt not make any ontological commitment". Indeed, this commandment brings us back to the initial question. Has the last word been spoken concerning the particle mass with the Higgs boson? Only in this case it was a valid argument that the absence of technical naturalness is an expression of a final, scale-crossing theory [^12] Already in 1979, Iliopoulos remarked (p. 89): ”Several people believe, and I share this view, that the Higgs scheme is a convenient parametrization of our ignorance concerning the dynamics of spontaneous symmetry breaking, and elementary scalar particles do not exist”. However, the detection of the Higgs particle is likely to refute this radical conclusion drawn from the quadratic UV sensitivity of scalar particles.. Koren (2020) concedes that a fundamental Higgs particle could indeed be the way the universe works. Nevertheless, she expects the Higgs mass to be a parameter resulting from input values of a UV theory in low-energy theory, even if it is unknown whether the Higgs mass ultimately emerges as part of a larger multiplet, as a bound state of fermions, or as an excitation in string theory. However, as we have already noted, those theories beyond the standard model, especially string theory, remain experimentally unconfirmed and pure speculation. Thus, it is hoped that the research program of BSM physics will ultimately lead to the desired situation, allowing for actual ontological statements about the smallest building blocks of matter for the first time.
However, how can the presumably fundamental nature of the Higgs field in the universe be reconciled with the fact that the standard model as a whole must be assigned to the class of EFTs, making final ontological conclusions premature? An answer to this question is provided by the approach of Selective Realism, formulated by Worrall (1989) and Kitcher (1993), among others. As a subcategory of Scientific Realism, it aims to label a thoughtful selection of individual theoretical assumptions as true or at least approximately true. The argument goes as follows: Having learned from the history of science, one should be cautious of untimely euphoria and not regard the currently best theories as truthful descriptions of reality. This principle is found under the term of pessimistic meta-induction according to Laudan (1981). However, the selective realist concedes that elements can be extracted from these theories that are likely to survive theory changes unscathed and will also be a central part of (all) improved theories in the future. They are, in a sense, the fixed points in the dynamic system of changing theories. Thus, the No-Miracle Argument is taken into account, suggesting that the outstanding empirical adequacy of the theory regarding the selected aspect would be inexplicable if it cannot be attributed at least approximately true. As Musgrave points out (1988), this argument can be considered the strongest support for all varieties of realism. However, given the evident limitations of the standard model to specific energy scales (EFT), individual realism that regards the entire model as (at least approximately) true is unjustified. This does not even require the argument of pessimistic meta-induction - it was, by construction, a provisional model that will one day be discarded. Whether the future replacement of the theory as a whole is known in advance or not, a selective realist avoids Laudan’s pessimism and argues that realistically interpretable aspects were causally linked to the temporary empirical adequacy of the theory. He attributes lasting value to them for theory-building and sees them as guarantors of success for future theories.
If an empirically successful aspect breaks the strong autonomy of scales, as in the case of the Higgs mass, the instrumentalist approach via EFTs, to which no even approximate truth can be ascribed, gradually comes to an end, if no experimental explanations for this behavior can be demonstrated. However, failure within the framework of such pragmatic theories need not be considered a problem, as the term "naturalness problem" has always suggested: Respecting the success of the theoretical description of the world using Higgs bosons, one should instead attribute to them a fundamental truth and cease the strained attempt to artificially force a fundamental phenomenon into the scope of provisional instrumentalist theories. Violations of strong autonomy of scales may be used as a criterion to already employ Selective Realism in the context of EFTs. However, it is crucial that the assumption of a fundamental nature of the respective phenomenon is justifiable. After all considerations, there is indeed nothing against the notion that particle masses are generated by coupling to the Higgs field on all energy scales. It is plausible that future theories will employ the same mechanism. A similar scrutiny is withstood by the previously discussed butterfly effect (an immutable principle in complex systems, also in refined future theories), noncommutative geometry (fundamental structure of spacetime in theories of quantum gravity), and the cosmological constant. On the other hand, the discussed phenomena with restored strong autonomy of scales occur in limited energetic domains (such as meson physics or classical electrodynamics). In light of this reconsideration, one can agree with Giudice (2017) that naturalness can still serve as an essential guiding principle even in the post-naturalness era, as long as its scope is understood.
In contrast, for example, the fermions of the standard model can indeed be considered fundamental and irreducible, even though their quantum corrections to mass through chiral symmetries are natural. As long as there are no signs of substructures of these assumed point-like and elementary particles, Selective Realism can also be applied here [^13] To avoid misunderstandings: Of course, the OSR is a sort of Selective Realism in another sense of this term as well. The purpose of OSR in this article, however, should not be confused with the application of Selective Realism, rescueing the potential contradiction of ontological statements about objects in EFTs. The OSR was only used as a tool to determine the physical reality of bare mass and quantum corrections. This kind of Selective Realism states which kinds of entities really exist fundamentally (structures instead of fundamental objects). Selective Realism in its second meaning just states that at each point within the evolution of a theory series, a certain quotient of already realistic elements occurs in the corresponding theory, whereas the remainder can still be considered as instrumental or pragmatic, thus metaphorical elements of the theory - providing a compromise between Scientific Realism and Instrumentalism in developed, but non-final theories.. Thus, violated strong autonomy of scales is not a necessary criterion for attributing (approximately) truth to theoretical assumptions, as I would like to explicitly emphasize. However, it can be understood as a sufficient criterion once the fundamental status of the phenomenon has been sufficiently made plausible.
As a perspective, we note that the lack of mathematical understanding of QFT carries the risk that its use in present-day particle physics could be fundamentally doomed. Grinbaum and Rivat (2020, p. 11) remark: "If the formalism of the new theory after a future theory change happens to be incommensurable with the mathematical structure of QFT, even the parts favored by selective realists will become void of ontological status". Even the Higgs boson with its UV-sensitive mass corrections is currently embedded in the formalism of QFT. We cannot exclude that a reformed version of QFT could shed new light on the problem of technical naturalness in its mathematical definition over $\Lambda^{2}$. The role of rigorous mathematics for an (approximately) true description of the world must therefore be thoroughly examined. Perhaps, to make ontological statements about the world, one cannot settle for effective or pragmatic mathematics. In this case, the phenomenological fundamentality of a phenomenon would not be sufficient to transition to Selective Realism. In other words, the question arises: To what extent must fundamental (non-pragmatic, not inconsistency-ignoring) mathematics and the description of fundamental phenomena go hand in hand for the theory to allow ontological statements? The background of this question is the idea that mathematics in foundational physics is not just a language for describing the world but is understood in this discipline as the world itself, an assemblage of mathematical structures. Ontology would then mean nothing less than writing down a set of mathematical equations and principles, as proposed by Tegmark (2008) among other authors. The mathematics used today to describe the (nevertheless, in our assumption, fundamental) Higgs boson, for example, could distort its true nature to such an extent that present-day physicists are not yet privy to any truth. Such distortion could occur, for example, in resolving the previously described Landau pole, which the gauge coupling to the hypercharge reaches at finite energies, prohibiting the limit $\Lambda\to\infty$. The residual jeopardy of the fundamental character of the Higgs boson in the current state of particle physics could be eliminated by the mathematics of the future, but also delay the applicability of Selective Realism.
5. Changes in the Aesthetic Canon of BSM Physics
Why is the critical stance towards technical naturalness or the autonomy of scales in general still a minority view? How is it possible that a concept, for which there is no evidence regarding its universal validity, has shaped a research program for decades, and even concrete experimental failures have not triggered a general rethinking in the vast majority of researchers? A philosophy of science explanation may be provided by McAllister’s model, which views scientific change as a change in aesthetic ideals.
If one were to summarize McAllister’s basic idea about scientific change in one sentence, the following quote would be suitable: "History of science consists essentially of the growth of truthlikeness or empirical success, and aesthetic canons rise or fall as they facilitate or impede that growth" (McAllister 1989, p. 46). At first, we summarize some basic ideas of the concept. For all properties $P_{i}$ of a theory $T$, a corresponding aesthetic criterion is assumed. Such a criterion is represented by two pieces of information: the property $P_{i}$ and a weighting $W_{P_{i}}$ with which the scientist attributes an aesthetic value to the property. It must be conceded that only a few properties $P_{i}$ are weighted differently from zero, namely those that can evoke an aesthetic sense in the scientist. Four classes of such aesthetic properties are intended to cover the majority of cases in McAllister’s considerations: symmetries of the theory, the invocation of a model, the visualizability or abstractness of the theory, and its metaphysical allegiance. The symmetry properties of physical theories have regularly been described as triggers of aesthetic feelings in their history (Penrose 1974, p. 271). However, the last three classes are somewhat further from the common concept of aesthetics.
Interpreting the underlying metaphysical assumptions as aesthetic properties of a theory may be particularly surprising. McAllister writes that indeed aesthetic properties are usually interpreted rather as a subset of metaphysical preferences (1996, p. 55). But on the one hand, the selection process of metaphysical principles is the same as for the selection of aesthetic symmetries or the degree of visualizability of a theory to be used. The scientific community selects those properties in each case that produce the greatest empirical success. The selection of metaphysical foundations becomes a category of aesthetic properties, not vice versa. On the other hand, violations of these chosen principles by the scientist cause (aesthetic) displeasure, analogous to the aversion to properties considered ugly or too abstract. In fact, in all cases, the acceptance of the theory is jeopardized by the same adherence to habits or (inter-)subjective assumptions. The inertia to abandon these as a scientific community corresponds to the irrationality highlighted by Kuhn on the eve of the revolution. For example, McAllister cites Einstein’s well-known aversion to quantum mechanics, which was based on the stoic retention of determinism as a metaphysical principle equally valid in the microcosm.
Let’s illustrate the four categories with aspects from the standard model. The presence of the aesthetic class of symmetries is hardly as evident in any research program as in particle physics: the gauge symmetries $SU(3)_{C}\times SU(2)_{L}\times U(1)_{Y}$, characterizing the three interactions of the SM through Lie groups, are just one example. Chiral symmetries or the hadron classification according to Gell-Mann (1961) extend this list and increase the aesthetic content of the theory. The invocation of a model occurs, for example, in the renormalization group flow, which is also successfully applied in statistical physics (Wilson 1975) or in the spontaneous symmetry breaking of a potential: a phenomenon that also occurs, for instance, in superconductivity. Abstract concepts of the SM can be visualized in everyday language, as shown by the concept of color charge in quantum chromodynamics. Finally, metaphysical principles can include the principle of naturalness, but also its concretization in the belief in the autonomy of scales. The hope for a grand unified theory, in which the interactions of the SM can be unified in a single large Lie group, can also be considered such a metaphysical principle.
In the extended sense, all these mentioned aspects can be understood as indicators of beauty. In McAllister’s conception, normal science lies in the productive coexistence of truth and beauty indicators as guiding principles in theory-building. There, one avoids the dilemma of having to choose between the aesthetic appeal of some theories and the observational success of others (McAllister 1989, p. 42). He describes the concept of "aesthetic induction", which transfers the aesthetic features of earlier theories to the present and assumes that they still withstand the empirical test. If this induction fails because, within the current aesthetic conventions, the hierarchically superior truth indicators do not resonate, the joint triumph of beauty and truth ends - a temporary success of artificially correlated indicators - and anomalies emerge (McAllister 1989, p. 43). Kuhnian irrationality now manifests itself in the aesthetic conservatism of some scientists who continue to evaluate theories guided by beauty indicators, regardless of their lower empirical success.
The following excerpt has a notable congruence with the adherence to technical naturalness: "As empirical progress is achieved, the prevailing indicators of beauty will be first strengthened, later undermined and ultimately destroyed by rising tensions between them and indicators of truth" (McAllister 1989, p. 46). In the success story of technical naturalness, we have mentioned three phenomena whose physical explanation initially reinforced the conviction of (strong) scale autonomy: beauty and truth indicators went hand in hand. Even today, these examples are repeatedly cited by proponents of strong scale autonomy like Giudice (2017). The turning point in this success story lies in the improper extrapolation of the principle to phenomena outside the scope of effective theories: If the metaphysical principle of decoupling distant energy scales is used in effective theories, it is repeatedly confirmed. When applied to (allegedly) fundamental phenomena that are valid on all energy and length scales, it fails. Any attempts to restore the scale autonomy of the Higgs mass through new physics increase the tension between the beauty indicator and the truth indicators since no such new physics has been empirically demonstrated. The universal adherence to effective field theories (EFTs) by a conservative fraction is replaced in this scenario by the distinction between effective and fundamental theories, each following its own metatheoretical principles.
In this debate, it becomes evident that Kuhn’s notion does not provide a suitable philosophy of science framework for the principle of naturalness as we will illustrate below [^14] In addition, also the Kuhnian incommensurability after paradigm shift is contradictory to the justified option that concepts like the Higgs boson, in the light of Selective Realism, will survive any future revolution.. McAllister’s model confirms that the transition to a new conception of theories in fundamental physics can often be considered in a rational discourse. Moreover, abandoning technical and ’t Hooft’s naturalness would not trigger a revolution. It would undoubtedly streamline the BSM research program, likely freeing it from pseudo-problems. It represents a return to the earlier, fruitful approach of physics, which seeks to resolve solely empirical or mathematical inconsistencies. This standpoint is one core message of Hossenfelder (2018), who already mentions a connection to McAllister without further elaborating. It recalls Einstein’s return to the principles of early classical physics, as described in (McAllister 1996) - although in this case, there were indeed inconsistencies between Newtonian mechanics and Maxwellian electrodynamics. Nevertheless, it reveals, as in BSM physics, that a conservatism of this particular kind, reminiscent of the successful foundations of earlier physics, has the potential for progressive problem shifts. Thus, a conservative approach does not necessarily entail an irrational adherence to outdated principles, as this assumption overlooks the nonlinear course of physics history with its loops and dead ends.
6. Conclusion
What impact does the discussion about the end of effective theories have on the near future of BSM physics? Regarding the forthcoming use of EFTs, we can only align ourself with the pragmatism of Georgi (1993, p. 215) writing: "In addition to being a great convenience, effective field theory allows us to ask all the really scientific questions that we want to ask without committing ourselves to a picture of what happens at arbitrarily high energy". Indeed, the broken autonomy of scales in the correction to the Higgs mass should be understood as a failure of the EFT for this specific scenario. Nevertheless, it can be managed with the choice of the bare Higgs mass, treating the emergence of UV sensitivity as a temporary mathematical artifact that disappears in renormalization, like all other significant corrections. The use of EFTs, limiting to specific energetic domains, remains a central prerequisite for a thriving particle physics in the future.
Nevertheless, the possibility of effective theories failing must be clearly recognised as an option when it comes to the dogmatically advocated technical naturalness and its limits. In the context of the presented arguments, the problem of the lacking technical naturalness of the Higgs mass would have to be treated as a pseudo-problem. But in the presented interpretation, it reveals itself as a glimmer of hope: A hint at the fundamental nature of the affected phenomenon, which to reveal is the long-term mission of fundamental physics. We believe that as long as one only dealt with the artificial restoration of strong autonomy of scales in the corrections to the Higgs mass through theories without empirical evidence, there was a debilitating element in the BSM research program. A reinterpretation of the broken strong autonomy of scales, as carried out in this work, denies it the status of a problem while simultaneously enabling a progressive shift. In this shift, the limit of the applicability of EFTs is accepted, legitimizing greater openness in theory-building and evaluation without attacking the general approach of BSM physics through effective theories, as described by Georgi. The article merely criticizes the dogmatic adherence to principles for which there are sufficient reasons to release them.
Acknowledgements
We thank Ulrich Krohs, Raimar Wulkenhaar, Michael te Vrugt und Florian Fabry for valuable discussions and all members of the colloquium for the discussion of publications of the Philosophisches Seminar (University of Münster) for a critical review.